OpenStudy (anonymous):

solving quad. equ. by factoring 3ksquare+72=33k

OpenStudy (anonymous):

\[3k^2+72=33k\] \[3k^2-33k+72=0\] \[3(k^2-11k+24)=0\] \[3(k-8)(k-3)=0\] \[k-3=0\] \[k=3\] \[k-8=0\] \[k=8\]